OpenStudy (datanewb):

Help getting started with the indefinite integral \[\int \frac{2 \, dx}{\sqrt{16 - x^2}} =\]

OpenStudy (anonymous):

i think this is the derivative of \(\arcsin(x)\) or at least it is close

OpenStudy (anonymous):

if you rewrite it as \[\frac{2}{4\sqrt{1-\left(\frac{x}{2}\right)^2}}\] then it is exactly that, except for the constant

OpenStudy (datanewb):

ahh, it's a trig identity... I think I can get it from here... I see a 1-cos^2 x = sin^2 buried in the denominator.. I can get it from there I think thank you!

OpenStudy (anonymous):

yw